OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + y = 6 6x + 2y = 8 zero 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

the equations(1) times 2.....you will know how to choice options

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

solve for x in any of the two equations then what you get apply it to the other equation

OpenStudy (anonymous):

y=6-3x

OpenStudy (anonymous):

6x+2(6-3x)=8

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

do you plus 2?

OpenStudy (anonymous):

6x+12-6x=8

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

you are left with 12=8

OpenStudy (anonymous):

thats not an answer though

OpenStudy (anonymous):

read your answers and see the end result.

OpenStudy (anonymous):

is it There is no x value as there is no solution to the system?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

is it There is no x value as there is no solution to the system?

OpenStudy (anonymous):

thats not an answer though

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

oh

OpenStudy (unklerhaukus):

parallel lines

OpenStudy (ikram002p):

@UnkleRhaukus check last number after = they are not parallel but mm they are confusing the solution lol

OpenStudy (unklerhaukus):

3x + y = 6 (i) y = -3x + 6 slope is -3, y intercept is 6 6x + 2y = 8 (ii) 2y = -6x + 8 y = -3x + 4 slope is -3, y intercept is 4 straight lines with the same slope but different y intercept are parallel. There is no point that can appear on both lines.

OpenStudy (anonymous):

Answer: There is no x value as there is no solution to the system.